# VisJudge-Bench: Aesthetics and Quality Assessment of Visualizations

### v1

@@ 5.2.1 Can MLLMs Assess Visualization Aesthetics and Quality Like Humans?
Metric Model Overall Fidelity Expressiveness Aesthetics
Readability Insight Design Style Composition Color
MAE (↓) Claude-3.5-Sonnet 0.823 0.977 0.902 1.152 0.782 0.939 0.862
Claude-4-Sonnet 0.618 0.839 0.757 0.83 0.678 0.733 0.785
Gemini-2.0-Flash 0.68 0.828 0.91 0.818 0.637 0.728 0.798
Gemini-2.5-Pro 0.661 1.241 0.944 0.898 0.839 0.918 0.98
GPT-4o 0.609 0.986 0.804 0.742 0.608 0.694 0.657
GPT-5 0.551 0.861 0.78 0.776 0.648 0.698 0.682
Qwen2.5-VL-7B 1.048 1.169 1.294 0.857 0.755 0.812 0.772
VisJudge 0.442 0.662 0.649 0.679 0.581 0.546 0.604
MSE (↓) Claude-3.5-Sonnet 1.006 1.573 1.303 1.982 0.99 1.463 1.198
Claude-4-Sonnet 0.596 1.18 0.974 1.142 0.771 0.932 1.037
Gemini-2.0-Flash 0.716 1.18 1.323 1.114 0.671 0.922 1.057
Gemini-2.5-Pro 0.674 2.287 1.477 1.36 1.108 1.322 1.46
GPT-4o 0.575 1.557 1.06 0.918 0.625 0.821 0.729
GPT-5 0.484 1.214 0.988 0.966 0.719 0.859 0.81
Qwen2.5-VL-7B 1.502 2.047 2.409 1.176 0.937 1.091 0.996
VisJudge 0.306 0.751 0.693 0.762 0.545 0.498 0.578
Corr. (↑) Claude-3.5-Sonnet 0.395 0.325 0.491 0.366 0.456 0.137 0.259
Claude-4-Sonnet 0.47 0.392 0.548 0.453 0.422 0.164 0.228
Gemini-2.0-Flash 0.395 0.371 0.458 0.418 0.46 0.157 0.209
Gemini-2.5-Pro 0.266 0.18 0.379 0.357 0.447 0.194 0.208
GPT-4o 0.482 0.382 0.539 0.442 0.472 0.277 0.363
GPT-5 0.429 0.256 0.438 0.383 0.463 0.277 0.295
Qwen2.5-VL-7B 0.322 0.34 0.349 0.278 0.356 0.148 0.155
VisJudge 0.681 0.571 0.625 0.572 0.567 0.512 0.385
@@
Table 3 presents a comprehensive performance comparison of seven representative models including the latest GPT-5 across our “Fidelity, Expressiveness, and Aesthetics” evaluation framework, revealing significant capability differences and systematic limitations in current MLLMs for visualization aesthetics and quality assessment.
@@
Hierarchical Capability Structure. Current MLLMs exhibit a clear hierarchical performance structure across evaluation dimensions. Models perform relatively well on “Fidelity” dimensions, reflecting their fundamental capability in identifying obvious data errors (e.g., proportion distortions, baseline issues). Within “Expressiveness” dimensions, models show better performance on Insight Discovery (average MAE 0.87) compared to Semantic Readability (average MAE 0.91). Most prominently, all models struggle significantly with “Aesthetics” dimensions across all three aesthetic sub-dimensions, with average MAE around 0.76 and most correlations below 0.3 (except Design Style at 0.44), highlighting the inherent challenges of subjective aesthetic assessment, which often involves nuanced cultural context and abstract design principles that are difficult for current models to grasp.
@@
Model-Specific Evaluation Characteristics. Through fine-grained analysis, we identify distinct “evaluation personalities” across different models. GPT-5 demonstrates balanced performance across dimensions with consistently competitive scores, particularly excelling in overall accuracy; GPT-4o shows relative strength in Color Harmony assessment (MAE 0.657), reflecting sensitivity to color aesthetics; Claude-4-Sonnet excels in Semantic Readability evaluation (MAE 0.757), showing advantages in information communication assessment; while Gemini-2.0-Flash leads in Data Fidelity (MAE 0.828), indicating focus on data accuracy. These differentiated capability distributions validate VisJudge-Bench’s diagnostic value and provide guidance for practical model selection.
@@
Domain-Specific Fine-tuning Effectiveness. Our specialized VisJudge achieves superior performance across all core metrics, with an overall MAE of 0.442 and correlation of 0.681. Among commercial models, GPT-5 achieves the best MAE performance (0.551) while GPT-4o reaches the highest correlation (0.482). Compared to these strong baselines, VisJudge demonstrates substantial improvements: 19.8% MAE reduction over GPT-5 (from 0.551 to 0.442) and 41.3% correlation improvement over GPT-4o (from 0.482 to 0.681), demonstrating the substantial potential of domain-specific fine-tuning.
Figure 4: Distribution and bias analysis of MLLM scores. Score distribution density curves showing the rating patterns of different models compared to human experts on the 1-5 scale.
@@ 5.2.2 Do MLLMs Exhibit Human-like Scoring Behaviors?
To analyze systematic biases in model evaluation behavior, we examine score distribution patterns across different models. Figure 4 reveals significant bias issues in current MLLMs compared to human experts ($\mu=3.13$).
Systematic Biases in Current Models. Most models exhibit score inflation with rightward-shifted distributions. Qwen2.5-VL-7B and Claude-3.5-Sonnet show the most severe inflation ($\mu=3.89$ and $\mu=3.87$), while Gemini-2.0-Flash, GPT-4o, Claude-4-Sonnet, and GPT-5 demonstrate moderate inflation ($\mu=3.64$, $\mu=3.53$, $\mu=3.56$, and $\mu=3.36$ respectively). Notably, GPT-5 shows relatively better control compared to other inflated models. Conversely, Gemini-2.5-Pro exhibits overly conservative behavior ($\mu=3.02$). Additionally, models like Qwen2.5-VL-7B, Claude-3.5-Sonnet, and Gemini-2.0-Flash exhibit sharp peaks around 4.0, indicating excessive score concentration that limits discriminative capability.
@@ 5.2.3 How Does Visualization Complexity Affect Model Performance?
To understand model robustness across varying complexity, we analyze eight models on three visualization types: single visualizations, multiple visualizations, and dashboards. Figure 5 shows the main trends.
@@
Stability in Complex Scenarios. Baseline models show significant instability in complex scenarios. For dashboards, most baselines experience substantial correlation drops, with Claude-3.5-Sonnet and GPT-5 even showing negative correlations in Data Fidelity (-0.031 and -0.013), while VisJudge maintains consistency (0.224-0.482). Functional dimensions (Data Fidelity, Semantic Readability) remain stable across types, but aesthetic dimensions struggle with complex layouts, particularly Visual Composition in dashboards (most models <0.2). These findings highlight the critical importance of specialized training for robust visualization evaluation across diverse complexity levels.
@@ 5.2.4 How Do Model Evaluation Behaviors Differ in Practice?
To qualitatively analyze model evaluation behaviors, our case studies reveal two common biases: “score inflation” and “overly conservative” assessments.
@@
Figure 6 illustrates score inflation on low-quality visualizations. For a chaotic treemap (human rating: 1.67), baseline models give inflated scores. For instance, Qwen2.5-VL-7B (3.67) praises its “clear legend” while ignoring the confusing layout, and Claude-4-Sonnet (3.08) incorrectly highlights “excellent spatial organization”. In contrast, VisJudge’s score of 2.00 aligns with human judgment, correctly identifying the “chaotic layout” that impairs interpretation.
Conversely, Figure 7 highlights the overly conservative bias of Gemini-2.5-Pro. For a high-quality dashboard rated 4.17 by humans, Gemini-2.5-Pro gives a disproportionately low score of 2.94, focusing on a single data inconsistency while overlooking the chart’s overall effectiveness. Similarly, for another chart (human rating: 3.56), it scores only 2.33 due to the use of dual Y-axes. While other models like GPT-5 and Claude-4-Sonnet provide scores closer to human ratings, VisJudge also demonstrates more balanced evaluations (3.83 and 3.00, respectively). For more detailed case studies and complete model outputs, see Appendix B. Specifically, we provide high-score cases (Appendix B.1), medium-score cases (Appendix B.2), and low-score cases (Appendix B.3) demonstrating human-model alignment across different quality levels. Additionally, dimension-specific case studies (Appendix B.4) validate our evaluation criteria, while comprehensive model error analysis (Appendix B.5) reveals systematic failure patterns.
@@
5.2.5 How Does Training Data Scale Affect Model Performance?
@@
Predictable Scaling Laws. Model performance follows well-defined trends: human-model correlation shows logarithmic growth (R²=0.628) from 0.30 to 0.65 at 2,442 samples, while prediction errors exhibit exponential decay with MAE decreasing from 1.05 to 0.45 (R²=0.984) and MSE from 1.55 to 0.30 (R²=0.992). The 500-1,000 sample range provides the most efficient improvement, contributing 45% of total correlation gains, while beyond 1,000 samples, marginal returns diminish but remain valuable for continued enhancement.
@@ 6 Conclusion
This paper constructs VisJudge-Bench and fine-tunes VisJudge to validate the effectiveness of domain-specific training. Our research finds that existing MLLMs (including GPT-5) show significant gaps with human experts in visualization evaluation, exhibiting issues like scoring bias. VisJudge effectively mitigates these problems, achieving 19.8% MAE reduction and 58.7% correlation improvement over GPT-5. VisJudge-Bench provides a standardized evaluation platform for the community, while VisJudge’s success demonstrates that domain-specific training is a viable approach for improving MLLMs’ evaluation capabilities, supporting future work on finer evaluation and higher-quality visualization generation.
@@
The VisJudge-Bench framework presented in this work aims to improve multimodal large language models’ capabilities in visualization quality assessment and promote the development of automated visualization evaluation technology. We believe this work will not produce direct negative social impacts, but recognize that the framework should be used with caution and ethical oversight when applied to sensitive domains or potentially harmful models. Although VisJudge-Bench aims to objectively assess visualization quality, the base models it relies on (such as Qwen2.5-VL-7B) or the datasets used to construct the benchmark may inadvertently reflect biases. Future work could investigate the fairness implications of these evaluation features across different populations, cultural backgrounds, and visualization styles. We particularly focus on the following ethical considerations: (1) strict compliance with copyright and usage terms during data collection; (2) ensuring fair compensation and voluntary participation for expert annotators; (3) avoiding content that may reinforce stereotypes or biases in benchmark design; (4) open-source release aimed at promoting community development rather than commercial monopoly.
@@
Appendix A. Dataset Construction Details ........................................................................................................................................................................A
A.1  Data Collection Process ........................................................................................................................................................................A.1
A.2  Dataset Statistics and Distribution ........................................................................................................................................................................A.2
A.3  Expert Annotation Process ........................................................................................................................................................................A.3
@@
Appendix C. Evaluation Framework Details ........................................................................................................................................................................C
@@
Appendix A Dataset Construction Details
A.1 Data Collection Process
A.1.1 Web Crawling and Multi-Stage Filtering Pipeline
@@
A.2 Dataset Statistics and Distribution
A.2.1 Complete Dataset Statistics

### v2

@@ 5.2.1 Can MLLMs Assess Visualization Aesthetics and Quality Like Humans?
Table 3 presents a comprehensive performance comparison of 12 representative models including the latest GPT-5 across our evaluation framework, revealing significant capability differences and systematic limitations in current MLLMs for visualization assessment.
@@
Hierarchical Capability Structure. Current MLLMs exhibit a clear hierarchical performance structure across evaluation dimensions. Models perform relatively well on “Fidelity” dimensions (average MAE 1.006), reflecting their fundamental capability in identifying obvious data errors (e.g., proportion distortions, baseline issues). Within “Expressiveness” dimensions, models show better performance on Insight Discovery (average MAE 0.883) compared to Semantic Readability (average MAE 0.991). Most prominently, all models struggle significantly with “Aesthetics” dimensions across all three aesthetic sub-dimensions, with average MAE around 0.755 and correlations of 0.408, 0.177, and 0.220 for Design Style, Composition, and Color respectively, highlighting the inherent challenges of subjective aesthetic assessment, which often involves nuanced cultural context and abstract design principles that are difficult for current models to grasp.
@@
Model-Specific Evaluation Characteristics. Through fine-grained analysis, we identify distinct “evaluation personalities” across different models. GPT-5 demonstrates balanced performance across dimensions with consistently competitive scores, particularly excelling in overall accuracy; GPT-4o shows relative strength in Color Harmony assessment (MAE 0.657), reflecting sensitivity to color aesthetics; Claude-4-Sonnet excels in Semantic Readability evaluation (MAE 0.757), showing advantages in information communication assessment; while Gemini-2.0-Flash leads in Data Fidelity (MAE 0.829), indicating focus on data accuracy. Among open-source models, Qwen2.5-VL-72B-Instruct achieves the best overall performance (MAE 0.702, Corr. 0.440), approaching commercial model capabilities, while smaller models like Qwen2.5-VL-3B-Instruct and Llava-v1.6-mistral-7B show larger performance gaps. These differentiated capability distributions validate VisJudge-Bench’s diagnostic value and provide guidance for practical model selection.
@@
Domain-Specific Fine-tuning Effectiveness. Our specialized VisJudge achieves superior performance across all core metrics. The best-performing variant based on Qwen2.5-VL-7B-Instruct reaches an overall MAE of 0.421 and correlation of 0.687. Among commercial models, GPT-5 achieves the best MAE performance (0.553) while GPT-4o reaches the highest correlation (0.482). Compared to these strong baselines, VisJudge demonstrates substantial improvements: 23.9% MAE reduction over GPT-5 (from 0.553 to 0.421) and 42.5% correlation improvement over GPT-4o (from 0.482 to 0.687), demonstrating the substantial potential of domain-specific fine-tuning. Notably, domain-specific fine-tuning consistently reduces MAE across different backbone architectures, yielding 30–40% error reductions on all three open-source families. Qwen2.5-VL-3B-Instruct achieves a 40.2% reduction (from 0.821 to 0.491), InternVL3-8B 31.8% (0.793 to 0.541), and Llava-v1.6-mistral-7B 31.5% (0.724 to 0.496), demonstrating robust generalization of our training methodology across diverse model families. In subsequent analyses, VisJudge refers to this best-performing variant (Qwen2.5-VL-7B-Instruct based) unless otherwise specified.
@@ 5.2.2 Do MLLMs Exhibit Human-like Scoring Behaviors?
To analyze systematic biases in model evaluation behavior, we examine score distribution patterns across representative baseline models. Figure 4 reveals significant bias issues in current MLLMs compared to human experts ($\mu=3.13$).
Figure 4: Distribution and bias analysis of MLLM scores. Score distribution density curves showing the rating patterns of different models compared to human experts on the 1–5 scale.
Systematic Biases in Current Models. Most models exhibit score inflation with rightward-shifted distributions. Qwen2.5-VL-7B-Instruct and Claude-3.5-Sonnet show the most severe inflation ($\mu=3.89$ and $\mu=3.87$), while Gemini-2.0-Flash, GPT-4o, Claude-4-Sonnet, and GPT-5 demonstrate moderate inflation ($\mu=3.64$, $\mu=3.53$, $\mu=3.56$, and $\mu=3.36$ respectively). Notably, GPT-5 shows relatively better control compared to other inflated models. Conversely, Gemini-2.5-Pro exhibits overly conservative behavior ($\mu=3.02$). Additionally, models like Qwen2.5-VL-7B-Instruct, Claude-3.5-Sonnet, and Gemini-2.0-Flash exhibit sharp peaks around 4.0, indicating excessive score concentration that limits discriminative capability.
@@ 5.2.3 How Does Visualization Complexity Affect Model Performance?
To understand model robustness across varying complexity, we analyze representative baseline models on three visualization types: single visualizations, multiple visualizations, and dashboards. Figure 5 shows the main trends.
@@
Stability in Complex Scenarios. Baseline models show significant instability in complex scenarios. For dashboards, most baselines experience substantial correlation drops, with Claude-3.5-Sonnet and GPT-5 even showing negative correlations in Data Fidelity (-0.031 and -0.013), while VisJudge maintains consistency (0.224–0.482). Functional dimensions (Data Fidelity, Semantic Readability) remain stable across types, but aesthetic dimensions struggle with complex layouts, particularly Visual Composition in dashboards (most models <0.2). These findings highlight the critical importance of specialized training for robust visualization evaluation across diverse complexity levels. For detailed error analysis of multi-visualization and dashboard cases, including systematic bias patterns and failure modes across different models, see Appendix B.6.
@@ 5.2.4 How Do Model Evaluation Behaviors Differ in Practice?
To qualitatively analyze model evaluation behaviors, we conduct case studies on representative baseline models to reveal two common biases: “score inflation” and “overly conservative” assessments.
@@
Figure 6 illustrates score inflation on low-quality visualizations. For a chaotic treemap (human rating: 1.67), baseline models give inflated scores. For instance, Qwen2.5-VL-7B-Instruct (3.67) praises its “clear legend” while ignoring the confusing layout, and Claude-4-Sonnet (3.08) incorrectly highlights “excellent spatial organization”. In contrast, VisJudge’s score of 2.00 aligns with human judgment, correctly identifying the “chaotic layout” that impairs interpretation.
Conversely, Figure 7 highlights the overly conservative bias of Gemini-2.5-Pro. For a high-quality dashboard rated 4.17 by humans, Gemini-2.5-Pro gives a disproportionately low score of 2.94, focusing on a single data inconsistency while overlooking the chart’s overall effectiveness. Similarly, for another chart (human rating: 3.56), it scores only 2.33 due to the use of dual Y-axes. While other models like GPT-5 and Claude-4-Sonnet provide scores closer to human ratings, VisJudge also demonstrates more balanced evaluations (3.83 and 3.00, respectively). Additional case studies across different quality levels and comprehensive model error analysis are provided in Appendix B.
Table 4: MatPlotAgent quality improvement with different feedback models.
Generation Model Direct Decoding Qwen2.5-VL-7B Qwen2.5-VL-72B VisJudge
(Baseline) as Feedback as Feedback as Feedback
GPT-5 67.15 66.49 69.80 72.07
GPT-4o 62.71 58.49 63.53 71.62
Gemini-2.5-Pro 68.10 65.04 66.12 72.20
Claude-4-Sonnet 64.50 67.25 65.20 72.23
Claude-3.5-Sonnet 63.89 59.23 63.14 69.04
Qwen2.5-VL-72B-Instruct 61.18 58.70 60.30 67.32
Qwen2.5-VL-7B-Instruct 49.76 45.38 44.94 55.29
5.2.5 Can VisJudge Generalize to Real-world Applications?
To validate practical generalization, we integrated VisJudge into two real-world visualization systems with significant distribution shifts from our training data (detailed setup in Appendix E.1).
Visualization Generation. VisJudge provides feedback to MatPlotAgent (Yang et al., 2024) for iterative quality improvement. Table 7 shows consistent improvements across seven generation models (+6.07 points average). Notably, base models without domain fine-tuning often degrade performance (7B: -2.39, 72B: -0.61 average). This degradation stems from their systematic evaluation biases (Section 5.2): score inflation and poor discriminative capability lead to misleading feedback that misdirects generation models toward suboptimal outputs. In contrast, VisJudge’s accurate quality assessment consistently improves all generators. Even state-of-the-art models like GPT-5 and Claude-4-Sonnet achieve +4.92 and +7.73 improvements, demonstrating that domain-specific fine-tuning is essential for providing effective feedback.
Visualization Recommendation. VisJudge as a reward model in HAIChart (Xie et al., 2024) improves recommendation accuracy on VizML dataset: +4.40% (Data Queries Hit@1), +2.50% (Overall Hit@1), +5.30% (Overall Hit@3). While 7B and 72B base models show minimal to moderate improvements, VisJudge achieves consistently larger gains across all metrics, outperforming both base models and the baseline (detailed results in Appendix E.1). This validates VisJudge’s broad applicability across generation and recommendation tasks.
@@ 5.2.5 Can VisJudge Generalize to Real-world Applications?
5.3 How Does Training Data Scale Affect Model Performance?
@@
Predictable Scaling Laws. Model performance follows well-defined trends: human–model correlation shows logarithmic growth (R²=0.678) from 0.341 to 0.639 as the training data scale increases to 2,442 samples, while prediction errors exhibit exponential decay with MAE decreasing from 0.847 to 0.440 (R²=0.964) and MSE from 1.045 to 0.309 (R²=0.981). The 500–1,000 sample range provides the most efficient improvement, contributing 41.7% of total correlation gains, while beyond 1,000 samples, marginal returns diminish but remain valuable for continued enhancement.
@@ 6 Conclusion
This paper constructs VisJudge-Bench and fine-tunes VisJudge to validate the effectiveness of domain-specific training. Our research finds that existing MLLMs (including GPT-5) show significant gaps with human experts in visualization evaluation, exhibiting issues like scoring bias. VisJudge effectively mitigates these problems, achieving 23.9% MAE reduction and 60.5% correlation improvement over GPT-5. VisJudge-Bench provides a standardized evaluation platform for the community, while VisJudge’s success demonstrates that domain-specific training is a viable approach for improving MLLMs’ evaluation capabilities, supporting future work on finer evaluation and higher-quality visualization generation.
7 Limitations and Future Work
While VisJudge-Bench demonstrates significant value in visualization quality assessment, it has inherent limitations. The dataset primarily focuses on static visualizations with limited coverage of dynamic evolution and interactive exploration. Additionally, since samples are mainly collected from the web where raw data is often unavailable, our evaluation framework aligns with real-world human assessment by primarily relying on visual presentation. Future work includes expanding dynamic visualization samples with temporal evaluation dimensions, exploring channels to acquire samples with raw data for more comprehensive data fidelity verification, and investigating distribution-based evaluation methods to better capture the diversity of human preferences. To facilitate reproducible research and community development, we will publicly release VisJudge-Bench along with all annotations, quality control records, and the trained VisJudge model weights. We welcome community contributions to improve benchmark quality and coverage.
@@ 7 Limitations and Future Work
The VisJudge-Bench framework presented in this work aims to improve multimodal large language models’ capabilities in visualization quality assessment and promote the development of automated visualization evaluation technology. We believe this work will not produce direct negative social impacts, but recognize that the framework should be used with caution and ethical oversight when applied to sensitive domains or potentially harmful models. Although VisJudge-Bench aims to objectively assess visualization quality, the base models it relies on (such as Qwen2.5-VL-7B) or the datasets used to construct the benchmark may inadvertently reflect biases. Future work could investigate the fairness implications of these evaluation features across different populations, cultural backgrounds, and visualization styles. We particularly focus on the following ethical considerations: (1) strict compliance with copyright and usage terms during data collection; (2) ensuring fair compensation and voluntary participation for expert annotators; (3) avoiding content that may reinforce stereotypes or biases in benchmark design; (4) open-source release aimed at promoting community development rather than commercial monopoly. All research was conducted in strict compliance with ICLR ethics guidelines.
@@
Appendix A. Dataset Construction Details and Statistical Analysis ........................................................................................................................................................................A
A.1  Data Collection and Construction Pipeline ........................................................................................................................................................................A.1
A.2  Dataset Statistics and Distribution Analysis ........................................................................................................................................................................A.2
A.3  Annotation Process ........................................................................................................................................................................A.3
A.4  Annotation Quality Analysis ........................................................................................................................................................................A.4
@@
B.6  Error Analysis of Multi Visualizations and Dashboard ........................................................................................................................................................................B.6
Appendix C. Evaluation Framework and Criteria ........................................................................................................................................................................C
@@
Appendix E. Extended Experiments ........................................................................................................................................................................E
E.1  Generalization to Real-World Applications ........................................................................................................................................................................E.1
E.2  Cross-Architecture Generalization ........................................................................................................................................................................E.1.3
@@
Appendix A Dataset Construction Details and Statistical Analysis
A.1 Data Collection and Construction Pipeline
@@
A.2 Dataset Statistics and Distribution Analysis
A.2.1 Overall Dataset Statistics and Distribution